IS TROPOLONE PLANAR? THE ROLE OF SYMMETRY ON PROTON-TRANSFER DYNAMICS

\underline{L. A. BURNS}, D. MURDOCK, and P. H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520-8107.

The interrelated phenomena of proton transfer and hydrogen bonding play vital roles in many chemical and biochemical processes. A model compound for investigating these concepts is tropolone where a potential barrier of finite height hinders the symmetric transfer of a hydron between hydroxylic (proton-donating) and ketonic (proton-accepting) oxygen atoms that are attached to adjacent positions on a conjugated seven-membered ring. Tropolone is computationally tractable, yet retains sufficient complexity to exhibit the multimode dynamics that govern proton transfer in substantially larger systems. Ab initio calculations on the structure of ground-state tropolone (X̃^{1}A_{1}) have been undertaken with particular emphasis toward evaluating the validity of symmetry constraints often imposed on geometry optimizations. The molecule is found to converge to a slightly nonplanar configuration at levels of theory which incorporate substantial electron correlation. Although small, the resulting barrier to planarity imparts significant anharmonicity to the lowest-frequency (ν_{39}) normal mode, suggesting the importance of other out-of-plane vibrations toward the promotion and state-specificity of proton transfer. The implications of this nonplanarity for the potential energy landscape of tropolone, as well as for the description of attendant proton-transfer pathways, will be discussed within the framework of the encompassing G_{4} molecular symmetry group.